# Adaptive Parametric Activation

### v2

@@ 4 Experiment Setup
MiSLAS [111] R50 [28] 61.7 51.3 35.8 52.7
KCL [43] 61.8 49.4 30.9 51.5
TSC [53] 63.5 49.7 30.4 52.4
RIDE (3E)+CMO[68] 66.4 53.9 35.6 56.2
DOC [86] 65.1 52.8 34.2 55.0
CC-SAM [116] 61.4 49.5 37.1 52.4
Our Baseline SE-R50 [37] 66.2 53.1 37.1 56.0
@@
APA* + AGLU (ours) 68.3${}^{\textbf{+1.9}}$ 54.8${}^{\textbf{+0.9}}$ 39.4${}^{\textbf{+2.1}}$ 57.9${}^{\textbf{+1.7}}$
RIDE (4E) [95] X50 [100] 68.2 53.8 36.0 56.8
SSD [52] 66.8 53.1 35.4 56.0
BCL [117] 67.9 54.2 36.6 57.1
CNT  [67] 63.2 52.1 36.9 54.2
ALA  [110] 64.1 49.9 34.7 53.3
ResLT  [13] 63.6 55.7 38.9 56.1
ABC-Norm [36] 60.7 49.7 33.1 51.7
RIDE (3E)+CMO+CR [60] 67.3 54.6 38.4 57.4
LWS+ImbSAM [115] 63.2 53.7 38.3 55.3
Our Baseline SE-X50 [37] 67.9 53.0 37.7 56.7
@@
APA* + AGLU (ours) 69.8${}^{\textbf{+1.6}}$ 55.70.${}^{\text{0.0}}$ 41.1+2.${}^{\textbf{+2.2}}$ 59.1${}^{\textbf{+1.7}}$
Implementation Details We primarily use Squeeze and Excite [37] as our baseline with ResNet-32 [28] for CIFAR-LT, ResNet50 for iNaturalist, ResNet50 and ResNext50 [100] for ImageNet-LT and ResNet152 for Places-LT, which has been pretrained on ImageNet1K according to [58]. For LVIS, we use SE-Resnets with MaskRCNN [26], FPN [54], Normalised Mask [89], RFS sampler [22] and GOL loss [1] as a baseline. For V3Det, we use SE-ResNet50 with FasterRCNN [72], FPN and Wrapped Cauchy classifier [23]. All baselines use bag of tricks [109], and strong training techniques that we discuss and ablate in the Appendix. Our motivation for using bag of tricks is two-fold, first, it pushes the performance even further and secondly, it showcases the generalisability of our work.
For our attention models, we replace the Sigmoid with APA, and we further use LayerNorm [4] and attention dropout [32] with $p=0.1$, for all datasets except iNaturalist. We denote this configuration as APA* in our Tables and we ablate its components. For our AGLU models, we simply use it in-place of RELU.
5 Results
Table 3: (a) Results of APA* and AGLU for iNaturalist and Places-LT. (b) Results of APA* and AGLU for CIFAR100-LT with imbalance 10 and 100.
@@ 5 Results
DisAlign [108] 70.6 39.3
MisLAS [111] 71.6 40.4
LADE [34] 70.0 38.8
ALA [110] 70.7 40.1
TSC [53] 69.7 -
CNT [67] - 39.2
WD+MaxNorm [3] 70.2 -
DOC [86] 71.0 -
BCL[117] 71.8 -
ResLT [13] 70.5 39.8
IIF [2] - 40.2
ABC Norm [36] 71.4 -
LWS+ImbSAM[115] 71.1 -
CC-SAM[116] 70.9 40.6
AREA [10] 68.4 -
Our Baseline with SE 71.3 40.5
@@
BALMS [71] 63.0 50.8
RIDE (4E) [95] - 49.4
ACE (4E) [7] - 49.6
DiVE [29] 62.0 45.4
SSD [52] 62.3 46.0
MisLas [111] 63.2 47.0
HSC [91] 63.1 46.7
LADE [34] 61.7 45.4
ResLT (3E) [13] 63.7 49.7
TLC (4E) [50] - 49.8
TSC [53] 59.0 43.8
RIDE+CMO [68] 60.2 50.0
CC-SAM [116] - 50.8
RIDE+CMO+CR[60] 61.4 50.7
AREA [10] 60.8 48.9
Our Baseline with SE 65.2 50.9
@@
APA*+AGLU(ours) 66.8${}^{\textbf{+1.6}}$ 52.3${}^{\textbf{+1.4}}$
Long-tailed Classification Benchmark. We compare APA* against ensemble and fusion models [95, 13, 50, 7], margin adjustment [34, 108, 110], contrastive learning [91, 53, 43], knowledge transfer [67], knowledge distillation [29, 52], decoupled methods [111, 2, 108, 36], sharpness aware minimisation [115, 116, 60] and data augmentation [111, 68].
On ImageNet-LT, as shown in Table 2, our baseline models with bag of tricks reach the state-of-the-art (SOTA) for both SE-ResNet50 (R50) and SE-ResNeXt50 (X50). We want to point out that most performance comes from the bag of tricks and not the SE module, as shown, in detail in the Appendix.
Our APA* outperforms the SE-R50 baseline by $1.4$ percentage points (pp) on average, by $1.3$pp on frequent categories, $1.2$pp on medium and $2.2$pp on few classes. Most importantly, it increases the performance of both frequent and rare classes, which is a unique advantage compared to the previous works. Additionally, our APA* exceeds RIDE with 3 Experts (3E) and CMO [68] by $1.2$pp on average and by $3.7$pp on the rare classes using a single model. When AGLU is combined, it pushes the performance of APA* R50, by $0.5$pp on average, by $0.8$pp on the frequent, $0.5$pp on the medium and $0.1$pp on the few classes.
APA* with X50 increases the SE-X50 baseline by $1.7$pp on average, $1.0$pp the many classes, $1.4$pp the medium classes and $1.7$pp the few classes. Furthermore, it outperforms RIDE(3E)+CMO+CR [60] by $1.0$pp on average, $1.6$pp on frequent classes, $0.8$ on common and $1.0$pp on rare classes using a single model. When AGLU is combined to APA* X50, it adds $0.7$pp on average, $0.9$pp on the frequent categories, $0.3$pp on the medium, and $1.7$pp on the few classes.
Table 4: For all ablations we use ImageNet-LT. In (a) we compare adaptive activation to other learnable activation functions using SE-ResNet50. In (b) we show that adaptive activation and AGLU generalise to other attention mechanisms. In (c) we show that APA* can be combined with classifier learning techniques. In (d) we show that APA* works effectively with deeper ResNets. In (e) we show the components of APA*. In (f) we compare AGLU with other activation functions.
@@
PC-Softmax [34] 56.0 57.4
cRT [44] 55.6 56.4
Decoupled-DRW  [8] 55.1 56.6
BSCE [71] 56.0 57.0
@@
PReLU [27] 54.8
ELU [11] 52.6
Mish [64] 57.4
GELU [31] 57.5
SiLU [31] 57.1
@@
On iNaturalist18, Places-LT and CIFAR100-LT, our SE baselines with bag of tricks again reach the SOTA, as shown in Table 3-a and b.
Regarding, iNaturalist18, APA* improves the SE-baseline by $1.0$pp. It also outperforms BCL [117] by $0.5$pp, ResLT by $1.8$pp and LWS+ImbSAM [115] by $1.2$pp. AGLU further enhances the performance by a staggering $3.5$pp compared to SE-baseline, which is a significant increase.
On Places-LT, our APA* also increases the performance of the SE baseline by $0.8$pp. Moreover, it surpasses MisLAS [111] by $0.9$pp and CC-SAM [116] by $0.7$pp. When AGLU is combined with APA*, it further increases the performance by $0.7$pp reaching $42.0\%$.
As shown in Table 3-b, on CIFAR100-LT, APA* improves the performance by $0.5$pp and $1.0$pp compared to SE baseline using an imbalance factor of 10 and 100 respectively. Furthermore, compared to RIDE+CMO [68], CC-SAM [116] and RIDE+CMO+CR [60], APA* achieves $1.9$pp, $1.1$pp and $1.2$pp higher accuracy respectively, under an imbalance factor of 100 using a single model. Finally AGLU, further boosts the performance of APA* by $1.1$pp and $0.4$pp for imbalance factor of 10 and 100 respectively.
Activation Ablation Study. We compare APA, without the Dropout and Layernorm, against different activation functions such as the Sigmoid and the Gumbel [1] using the SE-ResNet50 baseline. Also, we implement Sigmoid and Gumbel variants with learnable temperature, to further understand their difference to our adaptive activation. As shown in Table 4-a, Sigmoid with temperature achieves slightly better performance for the rare classes, however it over-fits the frequent categories. In contrast, our APA achieves the best performance, increasing the overall performance by $0.3$pp, the frequent classes by $0.2$pp and the common categories by $0.4$pp compared to the second best Gumbel with temperature.
Generalisation to other Attention mechanisms. In Table 4-b, we show that APA can be combined with other attention mechanisms. Specifically, APA improves the performance of Spatial Attention by $0.4$pp and the Spatial-Channel Attention by $1.3$pp. AGLU further increases the performance of Spatial Attention by $1.2$pp and Spatial-Channel Attention by $0.2$pp.
Table 5: Comparisons on LVISv1.0 using MaskRCNN-FPN and 2x schedule.
@@
RFS  [22] R50 23.7 13.3 23.0 29.0 24.7
IIF [2] 26.3 18.6 25.2 30.8 25.8
Seesaw [89] 26.4 19.6 26.1 29.8 27.4
LOCE [20] 26.6 18.5 26.2 30.7 27.4
PCB+Seesaw [30] 27.2 19.0 27.1 30.9 28.1
ECM  [39] 27.4 19.7 27.0 31.1 27.9
GOL  [1] 27.7 21.4 27.7 30.4 27.5
ECM+GAP [107] 26.9 20.1 26.8 30.0 27.2
@@
GOL+AGLU(ours) APA*-R50 29.1${}^{\textbf{+0.9}}$ 21.6${}^{\textbf{+0.2}}$ 29.6${}^{\textbf{+0.7}}$ 31.7${}^{\textbf{+0.6}}$ 29.0${}^{\textbf{+0.9}}$
RFS [22] R101 [28] 27.0 16.8 26.5 32.0 27.3
NorCal [65] 27.3 20.8 26.5 31.0 28.1
Seesaw [89] 28.1 20.0 28.0 31.8 28.9
GOL [1] 29.0 22.8 29.0 31.7 29.2
ECM [39] 28.7 21.9 27.9 32.3 29.4
PCB + Seesaw[30] 28.8 22.6 28.3 32.0 29.9
ROG [107] 28.8 21.1 29.1 31.8 28.8
@@
GOL+AGLU(ours) APA*-R101 30.7${}^{\textbf{+1.0}}$ 23.6${}^{\textbf{+0.6}}$ 31.3${}^{\textbf{+1.4}}$ 33.1${}^{\textbf{+0.7}}$ 31.11.${}^{\textbf{1.1}}$
Combining APA* with Classifier Learning. APA* is an efficient module that can be easily combined with common classifier learning techniques such as margin adjustment [71], reweighting [8] and resampling [44]. As shown in Table 4-c, APA* consistently boosts the performance of all these methods.
Larger ResNets. We further train deeper models like ResNet-101 and ResNet-152 on ImageNet-LT and compare the SE and APA* attention methods. As Table 4-d shows, the APA* module enlarges the performance of all models consistently for both frequent, common and rare classes. Especially, APA* with ResNet101 promotes overall accuracy by $2.1$pp, boosting the rare categories by a significant $4.5$pp compared to SE-Resnet101.
Ablation of APA* components. APA* uses Channel attention, APA, Dropout and Layer-Norm and we show their effects in Table 4-e. The plain R50, without channel attention, achieves $55.0\%$ and SE increases its performance by $1.0$pp. APA increases the performance of the SE baseline by another $1.0$pp. Dropout increases the performance by $0.3$pp and LayerNorm enhances the performance by an additional $0.1$pp. In the Appendix, we show the full component ablation.
Comparison of AGLU. In Table 4-f, we compare AGLU to other commonly used activations functions, using APA* ResNet50 as backbone and switching the activation function of the intermediate layers. Our AGLU outperforms all the other methods and it is the best choice.
Long-Tailed Instance Segmentation. As the results suggest in Table 5, the SE-R50 backbone increases the overall mask performance compared to plain GOL-R50 by $0.5$pp but most improvement comes from the common and frequent categories while the rare categories are significantly reduced by $0.8$pp. In contrast, our APA* with AGLU-R50 improves the performance by $0.9$pp on average mask and bounding box precision, $1.0$pp on $AP_{r}$, $0.7$pp on $AP_{c}$ and $0.9$pp on $AP_{f}$ compared to SE-R50-GOL. Compared to GOL with SE-R101, APA* with AGLU also improves the performance by $1.0$pp on $AP$, $0.6$pp on $AP^{r}$, $1.4$pp on $AP^{c}$, $0.7$pp on $AP^{f}$ and $1.1$pp on $AP^{b}$. This highlights that our APA* and AGLU modules are robust for the rare classes and they outperform the previous SOTA in long-tailed instance segmentation.
@@
Qualitative Results. We use APA* and Imagenet-LT for our qualitative analysis. In Fig. 5-(a) we show that APA* increases the variance of channel attention for the most semantic layer by $0.04$ compared to the baseline, making it diverse and informative for all channels. In Fig. 5-(b), we show that APA* increases the entropy of the attention signal for the deeper layers by $0.4\sim 0.6$ compared to the baseline, showing that our module produces informative signals that effectively attend to the rare classes. In Fig.-c 5-top, we show that the baseline channel attention of the rare classes, produces all-pass filters in other words, attention signals that have small entropy, i.e., $E=0.018$ for the bee house and $E=0.098$ for the warthog respectively. This hinders rare class learning and it results in misclassification. In contrast, our method in Fig. 5-bottom produces informative channel attention signals that have larger entropy i.e. $E=0.46$ and $E=0.91$ for the bee house and warthog classes respectively, allowing the model to retrieve rare class features and to make correct predictions. Visualisations of the learned $\kappa$ and $\lambda$ parameters are provided in the Appendix.

### v3

@@ V Experiments in Long-tail learning
MiSLAS [zhong2021improving] R50 61.7 51.3 35.8 52.7
KCL [kang2021exploring] 61.8 49.4 30.9 51.5
TSC [li2022targeted] 63.5 49.7 30.4 52.4
RIDE (3E)+CMO[park2022majority] 66.4 53.9 35.6 56.2
DOC [wang2022towards] 65.1 52.8 34.2 55.0
CC-SAM [zhou2023class] 61.4 49.5 37.1 52.4
Our Baseline SE-R50 66.2 53.1 37.1 56.0
@@
[…]
RIDE (4E) [wang2021long] X50 68.2 53.8 36.0 56.8
SSD [li2021self] 66.8 53.1 35.4 56.0
BCL [zhu2022balanced] 67.9 54.2 36.6 57.1
CNT  [parisot2022long] 63.2 52.1 36.9 54.2
ALA  [zhao2022adaptive] 64.1 49.9 34.7 53.3
ResLT  [cui2022reslt] 63.6 55.7 38.9 56.1
ABC-Norm [hsu2023abc] 60.7 49.7 33.1 51.7
RIDE (3E)+CMO+CR [ma2023curvature] 67.3 54.6 38.4 57.4
LWS+ImbSAM [zhou2023imbsam] 63.2 53.7 38.3 55.3
Our Baseline SE-X50 67.9 53.0 37.7 56.7
@@
[…]
V-B Implementation Details
We primarily use Squeeze and Excite [hu2018squeeze] as our baseline with ResNet-32 [he2016deep] for CIFAR-LT, ResNet50 for iNaturalist, ResNet50 and ResNext50 [xie2017aggregated] for ImageNet-LT and ResNet152 for Places-LT, which has been pretrained on ImageNet1K according to [liu2019large]. For LVIS, we use SE-Resnets with MaskRCNN [he2017mask], FPN [lin2017feature], Normalised Mask [wang2021seesaw], RFS sampler [gupta2019lvis] and GOL loss [alexandridis2022long] as a baseline. All baselines use bag of tricks [zhang2021bag], and strong training techniques that we discuss and ablate in the Appendix. Our motivation for using bag of tricks is two-fold, first, it pushes the performance even further and secondly, it showcases the efficacy of our work.
For our attention models, we replace the Sigmoid with APA, and we further use LayerNorm [ba2016layer] and attention dropout [hinton2012improving] with $p=0.1$, for all datasets except iNaturalist. We denote this configuration as APA* in our Tables and we ablate its components. For our AGLU models, we simply use it in-place of RELU, or GELU. Following [he2015delving] we disable the weight decay of AGLU parameters, allowing the activation curve to reach the optimal shapes during training.
TABLE III: (a) Results of APA* and AGLU for iNaturalist and Places-LT. (b) Results of APA and AGLU for CIFAR100-LT with imbalance 10 and 100.
@@ V-B Implementation Details
DisAlign [zhang2021distribution] 70.6 39.3
MisLAS [zhong2021improving] 71.6 40.4
LADE [hong2021disentangling] 70.0 38.8
ALA [zhao2022adaptive] 70.7 40.1
TSC [li2022targeted] 69.7 -
CNT [parisot2022long] - 39.2
WD+MaxNorm [alshammari2022long] 70.2 -
DOC [wang2022towards] 71.0 -
BCL[zhu2022balanced] 71.8 -
ResLT [cui2022reslt] 70.5 39.8
IIF [alexandridis2023inverse] - 40.2
ABC Norm [hsu2023abc] 71.4 -
LWS+ImbSAM[zhou2023imbsam] 71.1 -
$C$-SAM[zhou2023class] 70.9 40.6
AREA [chen2023area] 68.4 -
Baseline with SE 71.3 40.5
@@
BALMS [Ren2020balms] 63.0 50.8
RIDE (4E) [wang2021long] - 49.4
ACE (4E) [cai2021ace] - 49.6
DiVE [he2021distilling] 62.0 45.4
SSD [li2021self] 62.3 46.0
MisLas [zhong2021improving] 63.2 47.0
HSC [wang2021contrastive] 63.1 46.7
LADE [hong2021disentangling] 61.7 45.4
ResLT (3E) [cui2022reslt] 63.7 49.7
TLC (4E) [li2022trustworthy] - 49.8
TSC [li2022targeted] 59.0 43.8
RIDE+CMO [park2022majority] 60.2 50.0
CC-SAM [zhou2023class] - 50.8
RIDE+CMO+CR[ma2023curvature] 61.4 50.7
AREA [chen2023area] 60.8 48.9
Baseline with SE 65.2 50.9
@@
[…]
V-C Long-tailed Classification Benchmark.
We compare APA* against ensemble and fusion models [wang2021long, cui2022reslt, li2022trustworthy, cai2021ace], margin adjustment [hong2021disentangling, zhang2021distribution, zhao2022adaptive], contrastive learning [wang2021contrastive, li2022targeted, kang2021exploring], knowledge transfer [parisot2022long], knowledge distillation [he2021distilling, li2021self], decoupled methods [zhong2021improving, alexandridis2023inverse, zhang2021distribution, hsu2023abc], sharpness aware minimisation [zhou2023imbsam, zhou2023class, ma2023curvature] and data augmentation [zhong2021improving, park2022majority].
On ImageNet-LT, as shown in Table II, our baseline models with bag of tricks reach the state-of-the-art (SOTA) for both SE-ResNet50 (R50) and SE-ResNeXt50 (X50). We want to point out that most performance comes from the bag of tricks and not the SE module, as shown, in detail in the Appendix.
Our APA* outperforms the SE-R50 baseline by $1.4$ percentage points (pp) on average, by 1.31.3pp on frequent categories, 1.21.2pp on medium and 2.22.2pp on few classes. Most importantly, it increases the performance of both frequent and rare classes, which is a unique advantage compared to the previous works. Additionally, our APA* exceeds RIDE with 3 Experts (3E) and CMO [park2022majority] by 1.21.2pp on average and by 3.73.7pp on the rare classes using a single model. When AGLU is combined, it pushes the performance of APA* R50, by 0.50.5pp on average, by 0.80.8pp on the frequent, 0.50.5pp on the medium and 0.10.1pp on the few classes.
APA* with X50 increases the SE-X50 baseline by 1.71.7pp on average, 1.01.0pp the many classes, 1.41.4pp the medium classes and 1.71.7pp the few classes. Furthermore, it outperforms RIDE(3E)+CMO+CR [ma2023curvature] by 1.01.0pp on average, 1.61.6pp on frequent classes, $0.8$ on common and 1.01.0pp on rare classes using a single model. When AGLU is combined to APA* X50, it adds 0.70.7pp on average, 0.90.9pp on the frequent categories, 0.30.3pp on the medium, and 1.71.7pp on the few classes.
TABLE IV: For all ablations we use ImageNet-LT. In (a) we compare adaptive activation to other learnable activation functions using SE-ResNet50. In (b) we show that adaptive activation and AGLU generalise to other attention mechanisms. In (c) we show that APA* can be combined with classifier learning techniques. In (d) we show that APA* works effectively with deeper ResNets. In (e) we show the components of APA*. In (f) we compare AGLU with other activation functions.
@@
PC-Softmax [hong2021disentangling] 56.0 57.4
cRT [kang2019decoupling] 55.6 56.4
Decoupled-DRW  [cao2019learning] 55.1 56.6
BSCE [Ren2020balms] 56.0 57.0
@@
PReLU [he2015delving] 54.8
ELU [clevert2015fast] 52.6
Mish [misra2019mish] 57.4
GELU [hendrycks2016gaussian] 57.5
SiLU [hendrycks2016gaussian] 57.1
@@
On iNaturalist18, Places-LT and CIFAR100-LT, our SE baselines with bag of tricks again reach the SOTA, as shown in Table III-a and b.
Regarding, iNaturalist18, APA* improves the SE-baseline by 1.01.0pp. It also outperforms BCL [zhu2022balanced] by 0.50.5pp, ResLT by 1.81.8pp and LWS+ImbSAM [zhou2023imbsam] by 1.21.2pp. AGLU further enhances the performance by a staggering 3.53.5pp compared to SE-baseline, which is a significant increase.
On Places-LT, our APA* also increases the performance of the SE baseline by 0.80.8pp. Moreover, it surpasses MisLAS [zhong2021improving] by 0.90.9pp and CC-SAM [zhou2023class] by 0.70.7pp. When AGLU is combined with APA*, it further increases the performance by 0.70.7pp reaching $42.0\%$.
As shown in Table III-b, on CIFAR100-LT, APA* improves the performance by 0.50.5pp and 1.01.0pp compared to SE baseline using an imbalance factor of 10 and 100 respectively. Furthermore, compared to RIDE+CMO [park2022majority], CC-SAM [zhou2023class] and RIDE+CMO+CR [ma2023curvature], APA* achieves 1.91.9pp, 1.11.1pp and 1.21.2pp higher accuracy respectively, under an imbalance factor of 100 using a single model. Finally AGLU, further boosts the performance of APA* by 1.11.1pp and 0.40.4pp for imbalance factor of 10 and 100 respectively.
Activation Ablation Study. We compare APA, without the Dropout and Layernorm, against different activation functions such as the Sigmoid and the Gumbel [alexandridis2022long] using the SE-ResNet50 baseline. Also, we implement Sigmoid and Gumbel variants with learnable temperature, to further understand their difference to our adaptive activation. As shown in Table IV-a, Sigmoid with temperature achieves slightly better performance for the rare classes, however it over-fits the frequent categories. In contrast, our APA achieves the best performance, increasing the overall performance by 0.30.3pp, the frequent classes by 0.20.2pp and the common categories by 0.40.4pp compared to the second best Gumbel with temperature.
Generalisation to other Attention mechanisms. In Table IV-b, we show that APA can be combined with other attention mechanisms. Specifically, APA improves the performance of Spatial Attention by 0.40.4pp and the Spatial-Channel Attention by 1.31.3pp. AGLU further increases the performance of Spatial Attention by 1.21.2pp and Spatial-Channel Attention by 0.20.2pp.
TABLE V: Comparisons on LVISv1.0 using MaskRCNN-FPN and 2x schedule.
@@
RFS  [gupta2019lvis] R50 23.7 13.3 23.0 29.0 24.7
IIF [alexandridis2023inverse] 26.3 18.6 25.2 30.8 25.8
Seesaw [wang2021seesaw] 26.4 19.6 26.1 29.8 27.4
LOCE [feng2021exploring] 26.6 18.5 26.2 30.7 27.4
PCB+Seesaw [he2022relieving] 27.2 19.0 27.1 30.9 28.1
ECM  [hyun2022long] 27.4 19.7 27.0 31.1 27.9
GOL  [alexandridis2022long] 27.7 21.4 27.7 30.4 27.5
ECM+GAP [zhang2023reconciling] 26.9 20.1 26.8 30.0 27.2
@@
[…]
RFS [gupta2019lvis] R101 [he2016deep] 27.0 16.8 26.5 32.0 27.3
NorCal [pan2021model] 27.3 20.8 26.5 31.0 28.1
Seesaw [wang2021seesaw] 28.1 20.0 28.0 31.8 28.9
GOL [alexandridis2022long] 29.0 22.8 29.0 31.7 29.2
ECM [hyun2022long] 28.7 21.9 27.9 32.3 29.4
PCB + Seesaw[he2022relieving] 28.8 22.6 28.3 32.0 29.9
ROG [zhang2023reconciling] 28.8 21.1 29.1 31.8 28.8
@@
[…]
Combining APA* with Classifier Learning. APA* is an efficient module that can be easily combined with common classifier learning techniques such as margin adjustment [Ren2020balms], reweighting [cao2019learning] and resampling [kang2019decoupling]. As shown in Table IV-c, APA* consistently boosts the performance of all these methods.
Larger ResNets. We further train deeper models like ResNet-101 and ResNet-152 on ImageNet-LT and compare the SE and APA* attention methods. As Table IV-d shows, the APA* module enlarges the performance of all models consistently for both frequent, common and rare classes. Especially, APA* with ResNet101 promotes overall accuracy by 2.12.1pp, boosting the rare categories by a significant 4.54.5pp compared to SE-Resnet101.
Ablation of APA* components. APA* uses Channel attention, APA, Dropout and Layer-Norm and we show their effects in Table IV-e. The plain R50, without channel attention, achieves $55.0\%$ and SE increases its performance by 1.01.0pp. APA increases the performance of the SE baseline by another 1.01.0pp. Dropout increases the performance by 0.30.3pp and LayerNorm enhances the performance by an additional 0.10.1pp. In the Appendix, we show the full component ablation.
Comparison of AGLU. In Table IV-f, we compare AGLU to other commonly used activations functions, using APA* ResNet50 as backbone and switching the activation function of the intermediate layers. Our AGLU outperforms all the other methods and it is the best choice.
V-D Long-Tailed Instance Segmentation.
As the results suggest in Table V, the SE-R50 backbone increases the overall mask performance compared to plain GOL-R50 by 0.50.5pp but most improvement comes from the common and frequent categories while the rare categories are significantly reduced by 0.80.8pp. In contrast, our APA* with AGLU-R50 improves the performance by 0.90.9pp on average mask and bounding box precision, 1.01.0pp on $AP_{r}$, 0.70.7pp on $AP_{c}$ and 0.90.9pp on $AP_{f}$ compared to SE-R50-GOL. Compared to GOL with SE-R101, APA* with AGLU also improves the performance by 1.01.0pp on $AP$, 0.60.6pp on $AP^{r}$, 1.41.4pp on $AP^{c}$, 0.70.7pp on $AP^{f}$ and 1.11.1pp on $AP^{b}$. This highlights that our APA* and AGLU modules are robust for the rare classes and they outperform the previous SOTA in long-tailed instance segmentation.
@@
We use APA* and Imagenet-LT for our qualitative analysis. In Fig. 6-(a) we show that APA* increases the variance of channel attention for the most semantic layer by $0.04$ compared to the baseline, making it diverse and informative for all channels. In Fig. 6-(b), we show that APA* increases the entropy of the attention signal for the deeper layers by $0.4\sim 0.6$ compared to the baseline, showing that our module produces informative signals that effectively attend to the rare classes. In Fig.-c 6-top, we show that the baseline channel attention of the rare classes, produces all-pass filters in other words, attention signals that have small entropy, i.e., $E=0.018$ for the bee house and $E=0.098$ for the warthog respectively. This hinders rare class learning and it results in misclassification. In contrast, our method in Fig. 6-bottom produces informative channel attention signals that have larger entropy i.e. $E=0.46$ and $E=0.91$ for the bee house and warthog classes respectively, allowing the model to retrieve rare class features and to make correct predictions. Visualisations of the learned $\kappa$ and $\lambda$ parameters are provided in the Appendix.
VI Results on other tasks
In this section, we comprehensively evaluate the generalisation ability of AGLU activation beyond long-tail learning. We test AGLU in balanced classification, object detection, large-scale supervised pretraining and finetuning, Multi-modal LLMs, image generation with diffusion transformer and pure language models. In every subsection, we introduce the model architecture, training details, dataset and discuss the results.
VI-A Image Classification
We perform experiments on ImageNet1K [deng2009imagenet], using ResNets [he2016deep] and ViT [dosovitskiy2021an] models. We train Resnets with channel attention [hu2018squeeze], and spatial attention [woo2018cbam] for 100 epochs using learning rate 0.256, batch size 256, weight decay 6e-6, mixup coefficient 0.2, AutoAugment [cubuk2019autoaugment], training crop size 224 and random-resized-crops. We adopt ViT-S and B variants using a convolutional patch embedding [xiao2021early], layer-scale [touvron2021going], cosine positional embeddings and average pooling instead of class token [beyer2022better]. We train these models for 300 epochs using learning rate 0.002, weight decay 0.05, mixup coefficient 0.2, cutmix coefficient 1.0, batch size 2048, RandAugment [cubuk2020randaugment], repeated augmentations, color jittering, training crop size 192, random-resized-crops. For ViT-B, we add stochastic depth [huang2016deep] 0.2 to prevent overfitting.
As Table VI-a shows, our method increases the accuracy for R50, SE-R50 and CBAM-R50, by 0.60.6pp, 1.21.2pp and 0.60.6pp respectively. Also, it increases accuracy by 0.90.9pp on SE-R101 and 0.50.5pp on SE-R152 showing that AGLU is robust for any model size. In Table VI-b we show that AGLU has competitive performance for ViT models, increasing by 0.20.2pp and 0.30.3pp the ViT-S and ViT-B models respectively.
TABLE VI: Results on ImageNet1K using ResNets in (a) and vision Transformers in (b).
Method top-1
ResNet50 [hu2018squeeze] 76.9
w/ AGLU 77.5
SE-ResNet50 [hu2018squeeze] 77.5
w/ APA*+AGLU 78.7
CBAM-ResNet50 [hu2018squeeze] 78.3
w/ APA*+AGLU 78.9
SE-ResNet101 79.4
w/ APA* + AGLU 80.3
SE-ResNet152 80.3
w/ APA* + AGLU 80.8
Method top-1
ViT-S 80.2
w/ AGLU 80.4
ViT-B 82.5
w/ AGLU 82.8
(a) Using ResNets
(b) Using ViT
TABLE VII: Dataset details of downstream tasks.
Dataset Train size Test size Number of Classes
ImageNet[deng2009imagenet] 1,281,167 50,000 1000
iNaturalist2017 [van2018inaturalist] 579,184 95,986 5,089
iNaturalist2018 [van2018inaturalist] 437,513 24,426 8,142
iNaturalist2021 [van2018inaturalist] 2,686,843 500,000 10,000
CIFAR10 [krizhevsky2009learning] 50,000 10,000 10
CIFAR100 [krizhevsky2009learning] 50,000 10,000 100
Places365 [zhou2014learning] 1,803,459 36,499 365
Training on ImageNet21K. We test AGLU under large scale training by performing ImageNet21k pretraining and finetuning. For this experiment we use a SE-ResNet50 and a simple ViT-B architecture that strictly follows [beyer2022better]. We train these models using batch size 2048, learning rate 0.004, Adam with momentum [kingma2014adam], weight decay 0.02, mixup coefficient 0.2, training crop size 224, RandAugment and random-resized-crop for 90 epochs. After pretraining, we perform finetuning on 7 downstream tasks, including finegrained classification, scene classification and object classification, using the datasets in Table VII.
In more detail, we finetune the models for 50 epochs with learning rate 0.04, SGD, no weight decay, batch size 512, mixup coefficient 0.2, cutmix coefficient 1.0, RandAugment, color-jittering, label-smoothing 0.1, train-crop 224 and random-resized-crops.
TABLE VIII: Comparative downstream results on 7 popular image classification datasets using 224 training and validation crops.
Method Imagenet1K iNaturalist 2017 iNaturalist 2018 iNaturalist 2021 CIFAR10 CIFAR100 Places365 Average
SE-Resnet50 80.5 67.3 73.4 82.1 96.0 86.4 54.3 77.1
w/ APA* + AGLU 81.5 69.7 76.4 83.9 97.6 87.6 54.7 78.8
ViT-B 83.9 72.1 78.9 84.7 99.1 93.5 56.9 81.3
w/ AGLU 84.0 72.2 79.3 85.1 99.2 93.6 57.1 81.5
The results are shown in Table VIII. Our activation significantly enhances the performance of SE-ResNet50 by 1.91.9pp macro-average. AGLU, also increases the performance of VIT-B by 0.20.2pp macro average, showing that AGLU is a good activation function for both Transformer and CNN architectures.
VI-B Experiments with Object detection.
We perform experiments on COCO [lin2014microsoft], Objects365v2[shao2019objects365] and the recently proposed V3Det[wang2023v3det], which is a challenging large scale detection dataset with 13K classes and 243K images. We report $AP^{b}$ and $AP^{m}$ for COCO and $AP^{b}$ for V3Det and Objects365v2, using primarily SE-ResNets backbones pretrained on ImageNet1K. We use the mmdetection framework and train all models using 2x schedule.